The vertex form of the equation of a parabola is y=(x+7)^2 + 13. What is the standard form of the equation
Answer by ewatrrr(24785)

 
Hi,
The vertex form of the equation of a parabola is
y=(x+7)^2 + 13
y = x^2 + 14x + 49 +13
y = x^2 + 14x + 62
Check by completing the square
y = ax^2 + bx + c ⇒ y = a(x -(-b/2a)^2) - a(-b/2a)^2 + c
y = (x+7)^2 -49 + 62
y = (x+7)^2 +13